# Finite-Time Behavior of Erlang-C Model: Mixing Time, Mean Queue Length and Tail Bounds

### v2

@@ C.2. Proof of Corollary 5
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]\leq\left(1+\chi(\pi_{n,t},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}.$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{4n^{1-2\alpha}}{7(1+\delta)}}+\left(10+\frac{2}{\delta}\right)n^{2\alpha-1}}\times e^{-\frac{x}{2(1+\delta)}}$
@@
where $L_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}L_{n}=1$.
@@
$\left|\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{\varepsilon x}{2(1+\delta)}}\right]-\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon x}{2(1+\delta)}}\right]\right|\leq\chi\left(\pi_{n,t},\nu_{n}\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon x}{(1+\delta)}}\right]}.$
@@
(159) $\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{\varepsilon(q-n)}{2(1+\delta)}}\right]\leq\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{2(1+\delta)}}\right]+\chi\left(\pi_{n,t},\nu_{n}\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{(1+\delta)}}\right]}\leq\left(1+\chi\left(\pi_{n,t},\nu_{n}\right)\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{(1+\delta)}}\right]}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon(q-n)>x\right]\leq\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{\varepsilon(q-n)}{2(1+\delta)}}\right]e^{-\frac{x}{2(1+\delta)}}\leq\left(1+\chi\left(\pi_{n,t},\nu_{n}\right)\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon(q-n)}{(1+\delta)}}\right]}e^{-\frac{x}{2(1+\delta)}}$
And so, by applying the mixing results in Theorem 3 and the MGF bounds in Lemma 3, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}\text{ when }\alpha\in[1,+\infty),$
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}\text{ when }\alpha\in(1/2,1),$
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{4n^{1-2\alpha}}{7(1+\delta)}}+\left(10+\frac{2}{\delta}\right)n^{2\alpha-1}}\times e^{-\frac{x}{2(1+\delta)}}\text{ when }\alpha\in(0,1/2).$

### v3

@@ C.2. Proof of Corollary 5
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]\leq\left(1+\chi(\pi_{n,t},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}.$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{0.5e^{-\frac{2n^{1-2\alpha}}{7}}+6n^{2\alpha-1}}\times\sqrt{ex}e^{-\frac{x}{2}}$
@@
$\left|\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{\varepsilon_{n}x}{2(1+\delta)}}\right]-\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon_{n}x}{2(1+\delta)}}\right]\right|\leq\chi\left(\pi_{n,t},\nu_{n}\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon_{n}x}{(1+\delta)}}\right]}.$
@@
(159) $\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{\varepsilon_{n}(q-n)}{2(1+\delta)}}\right]\leq\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon_{n}(q-n)}{2(1+\delta)}}\right]+\chi\left(\pi_{n,t},\nu_{n}\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon_{n}(q-n)}{(1+\delta)}}\right]}\leq\left(1+\chi\left(\pi_{n,t},\nu_{n}\right)\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon_{n}(q-n)}{(1+\delta)}}\right]}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}(q-n)>x\right]\leq\mathbb{E}_{\pi_{n,t}}\left[e^{\frac{\varepsilon_{n}(q-n)}{2(1+\delta)}}\right]e^{-\frac{x}{2(1+\delta)}}\leq\left(1+\chi\left(\pi_{n,t},\nu_{n}\right)\right)\sqrt{\mathbb{E}_{\nu_{n}}\left[e^{\frac{\varepsilon_{n}(q-n)}{(1+\delta)}}\right]}e^{-\frac{x}{2(1+\delta)}}.$
Moreover, observe that $e^{\frac{x\delta}{\delta+1}}\geq\frac{x\delta}{\delta+1}+1\Leftrightarrow\sqrt{1+\frac{1}{\delta}}e^{-\frac{x}{2(1+\delta)}}\leq\sqrt{ex}e^{-\frac{x}{2}}\,\forall x\in\mathbb{R}$. And so, by applying the mixing results in Theorem 3 and the MGF bounds in Lemma 3, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}\text{ when }\alpha\in[1,+\infty),$
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}\text{ when }\alpha\in(1/2,1).$
Now when $\alpha\in(0,1/2)$ and $\delta=1$, we have
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{0.5e^{-\frac{2n^{1-2\alpha}}{7}}+6n^{2\alpha-1}}\times\sqrt{ex}e^{-\frac{x}{2}}.$